(SHOW YOUR WORK!!!) A local retailer claims that the mean lifetime of its lithium batteries is normally distributed with a mean of 1400 hours. A homeowner randomly selects 25 of these batteries and finds the mean lifetime to be 1340 hours with a standard deviation of 45 hours. Test the manufacturer’s claim. At an ? = .1, test the retailer’s claim.

a.) State the null and alternative hypotheses.

b.) Verify that the requirements are met for conducting the hypothesis test.

c.) Conduct the test of hypothesis by constructing a 90% confidence interval.
Answer #1

a)
Below are the null and alternative Hypothesis,
Null Hypothesis, H0: μ = 1400
Alternative Hypothesis, Ha: μ ≠ 1400

b)
Yes, requirements to conduct hypothesis test are ment
1. data is randomly selected from a population
2. The population data is normally distributed

c)
sample mean, xbar = 1340
sample standard deviation, s = 45
sample size, n = 25
degrees of freedom, df = n - 1 = 24

Given CI level is 90%, hence α = 1 - 0.9 = 0.1
α/2 = 0.1/2 = 0.05, tc = t(α/2, df) = 1.711

CI = (xbar - tc * s/sqrt(n) , xbar + tc * s/sqrt(n))
CI = (1340 - 1.711 * 45/sqrt(25) , 1340 + 1.711 * 45/sqrt(25))
CI = (1324.6 , 1355.4)

As 1400 is not part of the above CI, reject the null hypothesis